OpenStudy (anonymous):

As a part of a contest, a store is giving a special prize if a contestant can roll 3 standard six-sided number cubes, and get the same number on each. What is the probability a contestant will be successful?

OpenStudy (wolf1728):

There are 216 total combinations (6*6*6) for rolling 3 dice. There are only 6 combinations in which the same number appears on all 3 dice. So the probability = 6 / 216 = 1 / 36

OpenStudy (anonymous):

Thank you :)

OpenStudy (wolf1728):

u r welcome